Five years ago I was 5 times older than my son. In 8 years time i will be 3 times older than my son. How old am i today?
Answer by Alan3354(69443)

Five years ago I was 5 times older than my son. In 8 years time i will be 3 times older than my son. How old am i today?
--------------
F = father's age 5 years ago
S = son's age 5 years ago
----
F = 6S
F + 13 = 4(S + 13)
-----
6S + 13 = 4S + 52
2S = 39
The son was 19.5 then, he's 24.5 now.
F = 6S = 117 then.
Father is 122 now.
===================
Let's hope the kid is adopted.
---------------
1 time older = 100% older = 2 times as old
2 times older = 200% older = 3 times as old
3 times older = 300% older = 4 times as old
4 times older = 400% older = 5 times as old
etc.